If the acceleration due to gravity is 98m/s^2 and the radius of the earth is 6400km, calculate a value for the mass of the earth.
1
Expert's answer
2021-02-10T13:13:42-0500

Acceleration due to gravity is given by

"g=Gm_e/R_e^2\\m_e=gR_e^2/G\\=\frac{(9.8)(6400\times10^3)^2}{6.67\times10^{-11}}=6\times10^{24}kg"